# DoFlow: Causal Generative Flows for Interventional and Counterfactual Time-Series Prediction

### v1

@@ Appendix C Data Synthesis
C.1 Tree
@@
    $X_{i,t}=f(X_{i,t-1},X_{\mathrm{pa}(i),t-1},U_{i,t})=\beta_{i}X_{i,t-1}+\sum_{j\in\mathrm{pa}(i)}\tilde{\beta}_{j}X_{j,t-1}+U_{i,t}/4.$ (45)
@@
    $X_{i,t}=f(X_{i,t-1},X_{\mathrm{pa}(i),t-1},U_{i,t})=\beta_{i}X_{i,t-1}\,(|U_{i,t}|+0.5)+\sum_{j\in\mathrm{pa}(i)}\tilde{\beta}_{j}X_{j,t-1}.$ (46)
The root node $X_{1}$ is initialized over the interval $0,t_{0}]$ using a Chain process:
$X_{1,t}=\beta_{1}X_{1,t-1}+A\sin\!\left(\tfrac{2\pi t}{P}+\phi\right)+U_{1,t}.$ (47)
The coefficients $\{\beta_{i}\}$ and $\{\tilde{\beta}_{i}\}$ are selected from predefined sets to ensure that the generated time series contain sufficient noise but do not explode over time. Each parameter is drawn uniformly from its set for every experimental trial, and results are reported as averages across trials.
C.2 Diamond
@@
    $X_{i,t}=f(X_{i,t-1},X_{\mathrm{pa}(i),t-1},U_{i,t})=\beta_{i}X_{i,t-1}+\sum_{j\in\mathrm{pa}(i)}\tilde{\beta}_{j}X_{j,t-1}+U_{i,t}.$ (48)
@@
    $X_{i,t}=f(X_{i,t-1},X_{\mathrm{pa}(i),t-1},U_{i,t})=\exp(\beta_{i}X_{i,t-1})\cdot\frac{1}{2+|U_{i,t}|}+\sum_{j\in\mathrm{pa}(i)}\tilde{\beta}_{j}X_{j,t-1}.$ (49)
C.3 Fully Connected Layer (FC-Layer)
@@
    $X_{i,t}=f(X_{i,t-1},X_{\mathrm{pa}(i),t-1},U_{i,t})=\beta_{i}X_{i,t-1}+\sum_{j\in\mathrm{pa}(i)}\tilde{\beta}_{j}X_{j,t-1}+U_{i,t}.$ (50)
@@
    $X_{i,t}=f(X_{i,t-1},X_{\mathrm{pa}(i),t-1},U_{i,t})=\sqrt{\frac{1}{2}\left|\sum_{j\in\mathrm{pa}(i)}\tilde{\beta}_{j}X_{j,t-1}\right|+|U_{i,t}|}+\beta_{i}X_{i,t-1}.$ (51)
C.4 Chain Linear (Chain)
@@
    $X_{i,t}=f(X_{i,t-1},X_{\mathrm{pa}(i),t-1},U_{i,t})=\beta_{i}X_{i,t-1}+\sum_{j\in\mathrm{pa}(i)}\tilde{\beta}_{j}X_{j,t-1}+U_{i,t}.$ (52)
@@
    $X_{i,t}=f(X_{i,t-1},X_{\mathrm{pa}(i),t-1},U_{i,t})=\sqrt{\frac{1}{2}\left|\sum_{j\in\mathrm{pa}(i)}\tilde{\beta}_{j}X_{j,t-1}\right|+|U_{i,t}|}+\beta_{i}X_{i,t-1}.$ (53)
C.5 Interventional and Counterfactual Simulations
Each time series window has length $T=120$. We set the context window to $\tau=90$ and the forecasting window to $T-\tau=30$. Interventions begin immediately after the context window $\{1,2,\dots,\tau\}$. To obtain system-level interventions, we manipulate the root node(s) $X_{1,t}$, (or $X_{1,t},X_{2,t},X_{3,t}$ in the FC-Layer graph), throughout the forecasting window $\{\tau+1,\dots,T\}$.
To ensure that the intervention produces a substantial deviation from the original observed future, we construct interventions by shifting the root node(s) from earlier periods into the forecasting window, thereby imposing patterns that differ markedly from the original evolution.
Since the underlying structural causal models are known, interventional simulation proceeds by first sampling the exogenous noises $\{U_{i,\tau+1,T}\}_{i=2}^{K}$. The intervened values are then obtained by applying $\mathrm{do}(X_{1,\tau+1:T}=\gamma_{1,\tau+1:T})$ and propagating forward with the sampled noises to generate the intervened trajectories $\{\tilde{X}_{i,\tau+1:T}\}_{i=2}^{K}$. For the FC-Layer graph, the intervened parental set includes $\{X_{i,t}\}_{i=1,2,3}$.
@@ C.5 Interventional and Counterfactual Simulations
C.6 Metrics
We evaluate model performance using both Root Mean Squared Error (RMSE) and Maximum Mean Discrepancy (MMD). Let the test batch size be $B=128$, the groundt-ruth value for node $i$ at time $t$ in batch $b$ be $x_{i,t}^{(b)}$, and the corresponding model prediction be $\hat{x}_{i,t}^{(b)}$.
@@ C.6 Metrics
We run each algorithm 50 times, each time using different batches with batch size $B=128$. The final reported RMSE is obtained by first averaging over all nodes $i=1,\dots,K$, then averaging across the $N=100$ realizations, and finally averaging over the 50 runs. The reported standard deviation (std) is calculated from the variation across the 50 runs.
@@
Appendix D Comparisons to BGM settings
Bijective Generation Mechanisms (BGM). The BGM framework [47] shows that if the true structural mechanism $f$ is in the BGM class (bijective/strictly monotone in the exogenous noise), then any learned mechanism $\hat{f}$ that (i) is also in the BGM class and (ii) matches the observed distribution—i.e., for the same parents value $X$ and action $A$,
$\small\hat{f}(X_{t-},X_{\text{pa},t-},U_{t})$ $\stackrel{{\scriptstyle d}}{{=}}\ f(X_{t-},X_{\text{pa},t-},U_{t})$
Relationships. Because of Proposition 4.3 (enabled by the CNF), DoFlow implements a bijective-in-noise mechanism and is thus comparable to BGM in the sense that $\hat{f}$ is bijective in $U_{t}$. However, they are not the same: BGM additionally requires observational distribution matching to obtain class-level identifiability, whereas the proofs of Proposition 4.3 and Corollary 4.5do not assume such matching.
Under (A1)–(A3), Proposition 4.3 establishes a continuously differentiable bijection $g:U_{t}\to Z_{t}$ (independent of $(X_{t-},X_{\text{pa},t-}$), so the induced mechanism $\hat{f}$ implemented by DoFlow is bijective in the exogenous noise, i.e., DoFlow lies in the BGM class. With the additional assumption (A4), we can therefore directly invoke the counterfactual recovery result of the BGM framework [47].
Appendix E More Experimental Results
E.1 Empirical Validation of Assumption (A3)
We test (A3): $p_{\theta}(Z_{t}\mid H_{t-1})=q(Z_{t})=N(Z_{t};0,1)$, using a two-sample Maximum Mean Discrepancy (MMD). Specifically, we test the joint factorization $p(Z_{t},H_{t-1})=q(Z_{t})p(H_{t-1})$ by computing the MMD between the empirical joint sample $\{(z_{t},h_{t-1})\}$ and a synthetic joint sample $\{(z_{t}^{\prime},h_{t-1})\}$, where $z_{t}^{\prime}\sim q=N(0,1)$.
We employ a product kernel $k((z,h),(z^{\prime},h^{\prime}))=k_{Z}(z,z^{\prime})\cdot k_{H}(h,h^{\prime})$, where both $k_{Z}(z,z^{\prime})$ and $k_{H}(h,h^{\prime})$ are radial basis function (RBF) mixtures:
$k_{Z}(z,z^{\prime})$ $=\exp(-\frac{\|z-z^{\prime}\|^{2}}{2\sigma_{Z}^{2}}),$ (56)
$k_{H}(h,h^{\prime})$ $=\exp(-\frac{\|h-h^{\prime}\|^{2}}{2\sigma_{H}^{2}}).$
To set the bandwidths $\{\sigma_{Z}\}$ and $\{\sigma_{H}\}$, we first compute the pooled pairwise distances on the union of the observed and synthetic sets (e.g., $\tilde{Z}=Z\cup Z^{\prime}$). We then calculate the median of these distances, and define the bandwidths as $\tfrac{1}{2}\,\text{median}$. The empirical MMD is given as:
$\widehat{\mathrm{MMD}}^{2}_{u}=\frac{1}{n(n-1)}\sum_{i\neq j}k(x_{i},x_{j})+\frac{1}{n(n-1)}\sum_{i\neq j}k(y_{i},y_{j})-\frac{2}{n(n-1)}\sum_{i\neq j}k(x_{i},y_{j}),$ (57)
with $x_{i}=(z_{i},h_{i})$, $y_{i}=(z_{i}^{\prime},h_{i})$.
For more reliable comparisons, we additionally sample two groundt-ruth sets, $\{(z_{t}^{\prime},h_{t-1})\}$ and $\{(z_{t}^{\prime\prime},h_{t-1})\}$, where both $z_{t}^{\prime}\sim\mathcal{N}(0,1)$ and $z_{t}^{\prime\prime}\sim\mathcal{N}(0,1)$. The test is conducted on the linear additive “Tree” DAG, performed separately for each dimension, and the aggregated results are reported in the table below.
Table 2: Comparison of DoFlow and groundt-ruth samples under the Tree DAG using per-dimension MMD. Reported are the mean, minimum, and maximum values across dimensions.
Mean Min Max
DoFlow $3.7\times 10^{-2}$ $5.0\times 10^{-3}$ $8.3\times 10^{-2}$
True $2.4\times 10^{-2}$ $4.7\times 10^{-3}$ $6.9\times 10^{-2}$
As shown in Table 2, DoFlow achieves small two-sample joint MMD values, which are comparable to those obtained from the groundt-ruth independent samples $\{(z_{t}^{\prime},h_{t-1})\}$ and $\{(z_{t}^{\prime\prime},h_{t-1})\}$. This provides empirical support for the validity of Assumption (A3).
E.2 Computational Costs

### v2

@@ Appendix D Data Synthesis and Metrics Definition
In the simulations for different graph structures, the root node $X_{1}$ (for Tree, Diamond, and Chain), and the root nodes $X_{1},X_{2},X_{3}$ (for FC-Layer) are initialized over the interval $0,t_{0}]$ using a Chain process:
$X_{i,t}=\beta_{1}X_{i,t-1}+A\sin\!\left(\tfrac{2\pi t}{P}+\phi\right)+U_{i,t}.$ (45)
Here, $i\in\{1\}$ for Tree, Diamond, and Chain, and $i\in\{1,2,3\}$ for FC-Layer. Besides, $A$, $P$, and $\phi_{i}$ denote the amplitude, period, and phase offset of the sinusoidal driver, and $U_{i,t}$ represents independent exogenous noise $N(0,1)$.
In the Tree, Diamond, and Chain structures, the hyperparameters for the root node $X_{1,t}$- namely $A$, $P$, and $\phi$, are set to $1.0$, 2020, and $\frac{2\pi}{9}$, respectively. For the FC-Layer, these are set to $1.5$, $15$, and $\frac{4\pi}{9}$, respectively.
D.1 Tree
@@
    $X_{i,t}=f(X_{i,t-1},X_{\mathrm{pa}(i),t-1},U_{i,t})=\beta_{i}X_{i,t-1}+\sum_{j\in\mathrm{pa}(i)}\tilde{\beta}_{i\leftarrow j}X_{j,t-1}+U_{i,t}/4.$ (46)
@@
    $X_{i,t}=f(X_{i,t-1},X_{\mathrm{pa}(i),t-1},U_{i,t})=\beta_{i}X_{i,t-1}\,(|U_{i,t}|+0.5)+\sum_{j\in\mathrm{pa}(i)}\tilde{\beta}_{i\leftarrow j}X_{j,t-1}.$ (47)
For the Tree graph, we set the self-lag coefficients $\beta_{i}\in\{0.5,0.4,0.3,0.2,0.1,0.2,0.2,0.2\}$ for $i=1,\dots,8$ (in order), and define edge-specific parent coefficients $\tilde{\beta}_{i\leftarrow j}$ only on the directed edges $1\!\to\!2$ (0.3), $1\!\to\!3$ (-0.3), $2\!\to\!4$ (0.3), $2\!\to\!5$ (-0.3), $3\!\to\!6$ (0.5), $3\!\to\!7$ (-0.5), and $7\!\to\!8$ (0.5), with $\tilde{\beta}_{i\leftarrow j}=0$ for all other pairs.
D.2 Diamond
@@
    $X_{i,t}=f(X_{i,t-1},X_{\mathrm{pa}(i),t-1},U_{i,t})=\beta_{i}X_{i,t-1}+\sum_{j\in\mathrm{pa}(i)}\tilde{\beta}_{i\leftarrow j}X_{j,t-1}+U_{i,t}.$ (48)
@@
    $X_{i,t}=f(X_{i,t-1},X_{\mathrm{pa}(i),t-1},U_{i,t})=\exp(\beta_{i}X_{i,t-1})\cdot\frac{1}{2+|U_{i,t}|}+\sum_{j\in\mathrm{pa}(i)}\tilde{\beta}_{i\leftarrow j}X_{j,t-1}.$ (49)
[…]
D.3 Fully Connected Layer (FC-Layer)
@@
    $X_{i,t}=f(X_{i,t-1},X_{\mathrm{pa}(i),t-1},U_{i,t})=\beta_{i}X_{i,t-1}+\sum_{j\in\mathrm{pa}(i)}\tilde{\beta}_{i\leftarrow j}X_{j,t-1}+U_{i,t}.$ (50)
@@
    $X_{i,t}$ $=f(X_{i,t-1},X_{\mathrm{pa}(i),t-1},U_{i,t})=\beta_{i}X_{i,t-1}+\Bigg(\sum_{j\in\mathrm{pa}(i)}\tilde{\beta}_{i\leftarrow j}\,X_{j,t-1}\Bigg)+$ (51)
    $\gamma\!\left(\exp\!\Big(\sum_{j\in\mathrm{pa}(i)}\tilde{\beta}_{i\leftarrow j}\,X_{j,t-1}-1\Big)-1\right)-\gamma\,\tanh\!\Big(\sum_{j\in\mathrm{pa}(i)}\tilde{\beta}_{i\leftarrow j}\,X_{j,t-1}\Big)+U_{i,t}.$
[…]
D.4 Chain Linear (Chain)
@@
    $X_{i,t}=f(X_{i,t-1},X_{\mathrm{pa}(i),t-1},U_{i,t})=\beta_{i}X_{i,t-1}+\sum_{j\in\mathrm{pa}(i)}\tilde{\beta}_{i\leftarrow j}X_{j,t-1}+U_{i,t}.$ (52)
@@
    $X_{i,t}=f(X_{i,t-1},X_{\mathrm{pa}(i),t-1},U_{i,t})=\bigl(\beta_{i}X_{i,t-1}+0.5\bigr)\,\lvert U_{i,t}\rvert+\tilde{\beta}_{i\leftarrow i-1}X_{i-1,t-1}+\tilde{\beta}_{i\leftarrow 1}\,X_{1,t-1}$ (53)
For the Chain structure, we set $\beta_{i}=0.6$ for all $i=1,\dots,50$, and draw independent Gaussian noises $U_{i,t}\sim\mathcal{N}(0,\sigma_{i}^{2})$ with $\sigma_{i}=0.2$. For each non-root node $i=2,\dots,50$, the immediate-predecessor coefficient is fixed as $\tilde{\beta}_{i\leftarrow i-1}=0.2$, and the skip connection from the root is fixed as $\tilde{\beta}_{i\leftarrow 1}=0.7$.
D.5 Interventional and Counterfactual Simulations
Each time series window has length $T=120$. We set the context window to $\tau=90$ and the forecasting window to $T-\tau=30$. Interventions begin immediately after the context window $\{1,2,\dots,\tau\}$.
To rigorously evaluate DoFlow under challenging causal interventions, we design interventions that alter the root node(s) of the causal DAG, since manipulating root causes propagates system-wide effects through all downstream variables. Each root node follows a cyclic structural equation (45).
To ensure that the intervention produces a fundamentally different dynamic pattern, we construct the interventional trajectory by phase-shifting the root cycle by half a period, i.e., $P/2$. This yields:
$\gamma_{i,t}=\beta_{1}\,\gamma_{i,t-1}+A\sin\!\left(\tfrac{2\pi(t+P/2)}{P}+\phi\right)+U_{i,t},\quad t\in\{\tau+1,\dots,T\}$ (54)
where $i\in\{1\}$ for Tree, Diamond, and Chain, and $i\in\{1,2,3\}$ for FC-Layer. This corresponds to a complete inversion of the cyclic trend for the root nodes of each graph structure (including a noise term to create different intervention schedules across runs). Consequently, the interventional trajectory exhibits an antiphase pattern relative to the original, creating the most challenging regime for causal forecasting and counterfactual recovery.
Below, we illustrate how to simulate the true interventional and counterfactual trajectories using the SCMs defined above. For clarity, we present the procedure for the Chain, Diamond, and Tree structures, where the intervened root node is $X_{1,t}$. The FC-Layer case follows identically, except that interventions are applied to root nodes $i\in\{1,2,3\}$ as defined in (54).
The interventional simulation proceeds by first sampling the exogenous noises $\{U_{i,\tau+1,T}\}_{i=2}^{K}$. The intervened values are then obtained by applying $\mathrm{do}(X_{1,\tau+1:T}=\gamma_{1,\tau+1:T})$ and propagating forward with the sampled noises to generate the intervened trajectories $\{\tilde{X}_{i,\tau+1:T}\}_{i=2}^{K}$. For the FC-Layer graph, the intervened parental set includes $\{X_{i,t}\}_{i=1,2,3}$.
@@ D.5 Interventional and Counterfactual Simulations
D.6 Metrics
We evaluate model performance using both Root Mean Squared Error (RMSE) and Maximum Mean Discrepancy (MMD). Let the test batch size be $B=128$, the ground-truth value for node $i$ at time $t$ in batch $b$ be $x_{i,t}^{(b)}$, and the corresponding model prediction be $\hat{x}_{i,t}^{(b)}$.
@@ D.6.1 RMSE
For observational and interventional testing, we evaluate each method over 50 test batches of batch size $B=128$. Within a batch, RMSE is computed by (i) averaging over nodes $i=1,\ldots,K$, then (ii) averaging over $N=100$ stochastic realizations, and finally (iii) averaging over the 50 batches. The reported standard deviation (std) reflects variation across the 50 batches.
For counterfactual testing, each method produces a single deterministic counterfactual trajectory for a given test instance, so $N=1$. We evaluate performance by averaging RMSE over counterfactual test batches (each with $B=128$), and report std across 50 batches.
D.6.2 MMD
@@
D.6.3 CRPS
For probabilistic forecasting, we evaluate the quality of the predictive distribution using the Continuous Ranked Probability Score (CRPS) [Matheson and Winkler, 1976]. For a univariate predictive CDF $F$ and a realized outcome $y\in\mathbb{R}$, the CRPS is defined as
$\mathrm{CRPS}(F,y):=\int_{-\infty}^{\infty}\Big(F(z)-\mathbf{1}\{z\geq y\}\Big)^{2}\,dz,$ (57)
where $\mathbf{1}\{\cdot\}$ is the indicator function.
In our experiments, $F$ is represented by an ensemble of samples $\{\hat{x}_{i,t}^{(b,n)}\}_{n=1}^{N}$ for node $i$, time $t$, and test sequence $b$, with realized value $x_{i,t}^{(b)}$. We use the standard ensemble approximation:
$\mathrm{CRPS}_{i,t}^{(b)}\approx\frac{1}{N}\sum_{n=1}^{N}\bigl|\hat{x}_{i,t}^{(b,n)}-x_{i,t}^{(b)}\bigr|-\frac{1}{2N^{2}}\sum_{n=1}^{N}\sum_{n^{\prime}=1}^{N}\bigl|\hat{x}_{i,t}^{(b,n)}-\hat{x}_{i,t}^{(b,n^{\prime})}\bigr|.$ (58)
We report CRPS averaged over all test sequences $b$, time steps $t\in\{\tau+1,\dots,T\}$, nodes $i=1,\dots,K$, and runs, analogous to the aggregation used for RMSE.
Notably, CRPS is applicable only to probabilistic models. Therefore, we report CRPS results only for DoFlow, DeepVAR, and MQF2.
Appendix E Model and Implementation Details
This section summarizes the main implementation and hyperparameter choices used in all experiments.
Common setup.
All models are implemented in PyTorch, with continuous normalizing flows (CNFs) integrated using torchdiffeq.odeint. Unless otherwise stated, we use: Adam optimizer with learning rate $10^{-3}$, batch size $128$, and ReLU activations in the CNF networks. For each node $i$ in the DAG, we instantiate for our algorithm: (i) an LSTM-based RNN encoder that takes as input the node value (and, when applicable, time features), and (ii) a CNF whose velocity field is parameterized by a 3-layer MLP with hidden width $64$. All time series are normalized per window using a mean–variance scaler over the context window.
E.1 Adapting observational baselines for interventional rollouts
For each baseline method, we enforce the causal factorization by training a separate model for each node $i$, using as input a rolling context window consisting of the node history and its parents’ histories, $\{X_{i,1:t-1},X_{\mathrm{pa}(i),1:t-1}\}$, and predicting the next-step target $X_{i,t}$. For architectures that natively support multi-horizon prediction, we use the same model class but set the prediction length to one step and apply it recursively.
To predict interventional future under an intervention schedule $\mathcal{I}$ with values $\{\gamma_{i,t}\}$, we perform multi-step rollout over the forecast window by iterating $t=\tau+1,\ldots,T$ and forecasting nodes in topological order: if $(i,t)\in\mathcal{I}$ we fix $\hat{x}_{i,t}=\gamma_{i,t}$; otherwise we predict $\hat{x}_{i,t}$ from the baseline model using the latest rolled-out histories. The rolled-out predictions (including fixed intervention values) are then fed back as inputs for subsequent time steps, ensuring that interventions propagate to downstream nodes through the DAG.
E.2 Synthetic SCM Experiments
For the synthetic DAG experiments (Tree, Diamond, Chain, and FC-Layer graphs), we first generate long multivariate time series using the SCMs in Appendix D. We then extract sliding windows for training and evaluation.
Each window has total length $T=120$, with a context window of $\tau=90$ steps and a forecasting horizon of $T{-}\tau=30$ steps. We construct windows with stride $1$ along the simulated sequence. The resulting windows are split along the time index into $80\%$ for training and $20\%$ for testing, using the first $0.8N$ time stamps for training and the remaining $0.2N$ for testing.
For all synthetic structures, we set up our algorithm:
  • Per-node RNN encoder: 3-layer LSTM with hidden size $15$.
  • Per-node CNF: MLP with hidden dimension $64$ and $3$ layers; the conditioning dimension is
    $\text{cond\_dim}_{i}=15+|\mathrm{pa}(i)|\cdot 15.$
We train DoFlow for $15$ epochs on the synthetic datasets. For interventional evaluation, we draw $100$ flow samples per window to plot the confidence interval as well as to calculate the metrics.
E.3 Hydropower Forecasting
For the hydropower case study, we use processed Argonne hydropower data at individual locations. From the raw time-stamped series, we extract windows of length
$\text{context\_length}=180,\quad\text{prediction\_length}=30,$
with stride $5$ along time. Thus each training and test window contains $180$ context time steps followed by a $30$-step forecasting horizon.
We incorporate the time features, including the time of day and calendar features, to obtain a time-feature dimension of $\mathrm{TF}=6$. We set up our algorithm:
  • Per-node RNN encoder: 3-layer LSTM with hidden size $15$.
  • Per-node CNF: MLP with hidden dimension $64$ and $3$ layers; the conditioning dimension is
    $\text{cond\_dim}_{i}=15+|\mathrm{pa}(i)|\cdot 15,$
including the node’s own hidden state, and the hidden states of its parents.
For training (performed offline and loaded from checkpoints in the main hydropower experiments), we use batch size $128$, Adam with learning rate $10^{-3}$, and $N_{\text{epochs}}=15$ with early stopping based on node-wise training loss.
E.4 Cancer Treatment Effect
For the cancer-treatment experiment, we use the public dataset from Bica et al. [2020a], preprocessed into patient-level sequences with actions and outcomes. We construct per-patient windows of length
$\text{context\_len}=55,\quad\text{pred\_len}=5,$
where the context contains the past tumor volumes and treatments, and the prediction horizon contains the next $5$ steps to be modeled under different treatment policies.
Here we distinguish between outcome and action histories via two separate RNNs in our algorithm:
  • Outcome RNN: 4-layer LSTM with hidden size $H=15$ and input size $1$ (normalized cancer volume).
  • Action RNN: 4-layer LSTM with hidden size $H=15$ and input size $D_{A}=4$ (chemo application, radio application, chemo dose, radio dose).
The CNF for the cancer outcome $Y_{t}$ is parameterized as a 3-layer MLP with hidden width $64$, with conditioning dimension
$\text{cond\_dim}=15+15,$
corresponding to the concatenation of the current outcome hidden state and action hidden state.
We construct a training set of patient windows via CancerTrainDataset with batch size $128$ and shuffle the patients at each epoch. We train the cancer-treatment model for $15$ epochs, using the same per-window mean–variance normalization as in the other experiments. All reported interventional and counterfactual metrics are computed by rolling out DoFlow over the $5$-step prediction horizon under alternative (simulated) treatment policies, as detailed in Section F.4.
Appendix F More Experimental Results
F.1 Computational Costs